### v1

# LLM-Text Watermarking based on Lagrange Interpolation

Abstract
The rapid advancement of LLMs (Large Language Models) has established them as a foundational technology for many AI- and ML-powered human–computer interactions. A critical challenge in this context is the attribution of LLM-generated text – either to the specific language model used or to the individual user who generated it. This is essential for combating misinformation, fake news, misinterpretation, and plagiarism. One of the key techniques for addressing this issue is watermarking.
This work presents a watermarking scheme for LLM-generated text based on Lagrange interpolation, which enables the recovery of a secret author identity even when the text has been heavily redacted by an adversary. The core idea is to embed a continuous sequence of points $(x,f(x))$ that lie on a single straight line. The $x$-coordinates are generated pseudorandomly using either an LFSR (when security is not a priority) or a cryptographically secure NFSR for high-security applications. The scheme’s efficiency and resilience to adversarial modifications are analysed. Experimental results show that the proposed method is highly effective, allowing the recovery of the author’s identity when as few as three points survive adversarial manipulation.
1 Introduction
Artificial Intelligence (AI) and Machine Learning (ML) have advanced significantly in recent years and have given rise to powerful models based on deep learning. In particular, Deep Neural Networks (DNNs) have enabled machines to perform complex cognitive tasks such as image recognition, speech processing, and natural language understanding with accuracy almost equivalent to human performance. A pivotal advancement in this domain is the emergence of Large Language Models (LLMs). They can seen as deep learning architectures trained on massive textual datasets, which can generate coherent, contextually relevant, and human-like text. LLMs operate by learning rich representations of language that capture its semantic, syntactic, and pragmatic patterns. These models, often based on the Transformer architecture, are pre-trained on diverse corpora and fine-tuned for specific tasks such as question answering, summarisation, translation, and dialogue. Their capabilities stem from their scale, both in terms of parameters and data, which allows them to generalise across a wide range of linguistic contexts.
Notable examples of LLMs include GPT-4 by OpenAI [16], which demonstrates strong reasoning and multitask capabilities; PaLM by Google [5], which emphasizes multilingual and multitask learning; and LLaMA by Meta [21], which explores high-performance language modeling with relatively modest computational resources. These models not only advance the state of natural language processing but also raise new challenges and opportunities for applications in domains such as content generation, code synthesis, and interactive AI systems.
The growing adoption of Large Language Models (LLMs) offers a wide array of practical benefits, including automated text correction, assistance with writing, and content refinement. However, alongside these advantages lies a more troubling aspect: the potential misuse of LLMs. One such concern arises when individuals present AI-generated content as their own, falsely claiming authorship. This challenge naturally raises an important question: how can we embed a watermark into LLM-generated text that allows a recipient to reliably trace its origin, whether to a specific user or to the LLM system itself?
The objective of such watermarking is to make it possible to attribute the text to its true source, thereby establishing authorship beyond reasonable doubt. This capability is especially critical in combating misinformation, propaganda, automated plagiarism, and the spread of fake news. In this paper, we take a step toward a simple yet effective watermarking method for LLM-generated text, aiming to create a foundation for trustworthy AI-generated communication.
2 Overview and Taxonomy of Watermarking in LLMs
The primary motivation behind watermarking has long been the need to verify that a document or digital artifact originates from its rightful source [8]. Classical watermarking enables creators to prove ownership of multimedia content beyond a reasonable doubt. In the era of generative AI and LLMs, this requirement is more urgent than ever, due to the growing difficulty in distinguishing between human and machine generated text. Watermarking for LLMs serves several essential purposes: (1) authenticating the origin of text, (2) preventing unauthorised reuse or plagiarism, (3) identifying misuse (e.g., misinformation, impersonation), and (4) Establishing accountability in high-stakes domains (e.g., education, journalism, code generation). Recent literature has explored a diverse range of strategies for embedding and detecting watermarks in LLM-generated content. These approaches can be organised into four broad categories:
Watermarking During Training
– In this approach, the watermark is embedded by modifying the model training data or objectives, so that identifiable patterns are learned and can be later detected. In the work [19], the authors introduced the notion of text radioactivity. It allows to detect whether an LLM has memorised radioactive (marked) data points. TextMarker described in [14] uses backdoor-based membership inference to embed private ownership information. The scheme [9] injects plausible yet fictitious knowledge into training corpora to create semantically valid but watermarkable patterns.
Watermarking at the Logit Level
– the output probability distribution (logits) is subtly perturbed before sampling tokens, embedding statistical patterns without altering the model weights. The work [12] splits the collection of tokens into green and red subsets. The green tokens encode watermarks while red ones mark fake ones. Wong et. al. [22] describes their logits-to-text watermarking, which jointly optimises encoder-decoder architecture for high-fidelity and high-detectability watermarks. The GumbelSoft watermarking described in [10] improves sampling diversity while maintaining robust detection by refining Gumbel-based logits control.
Watermarking During Token Sampling
– this approach injects watermarks during the token selection process, often by biasing token choices at generation time. In their work [12], Kirchenbauer et al. applies a stochastic green token sampling, which softly biases generation toward watermark-encoding tokens at each step. The work [23] proposes frequency-based watermarking that embeds periodic token patterns detectable via Fourier analysis. In their work [25], Zhu et al. discuss their dual-pattern embedding, which simultaneously encodes watermarks in both the logits and sampling distributions.
Post-Generation Detection and Attribution
– the following works focus on detecting watermark patterns or attributing text to specific models or users after generation. Niess et al. [15] proposes a method for generating watermarks that alters token probabilities using stylometry. More precisely, it uses acrostica and sensorimotor norms to LLMs. Zhao et al. [24] introduce their Unigram-Watermark, which a theoretical framework for watermarking generation and correctness of watermarking detection.
For a comprehensive survey of LLM watermarking techniques, readers are referred to the review by Liu et al. [13].
2.1 Contributions
The paper makes the following contributions:
  • $\bullet$
    We design a watermarking scheme that embeds points from a straight line $f(x)$ into LLM-generated text. The $x$-coordinates are generated pseudorandomly, and the watermark consists of the sequence of points $(x,f(x))$. To extract the watermark, the verifier must solve the Maximum Collinear Points problem. We show that several algorithms exist for this task, with the most efficient ones having quadratic complexity, i.e., ${\cal O}(N^{2})$, where $N$ is the number of points.
  • $\bullet$
    We propose a variant of the watermarking scheme that is immune to algebraic attacks. This version employs cryptographically secure generation of $x$-coordinates and dynamic partitioning of the token set into ${\cal V}_{0}$ and ${\cal V}_{1}$.
  • $\bullet$
    Our probabilistic analysis demonstrates that any adversarial modification of the points $(x,f(x))$ is likely to be detected, as the altered points are statistically scattered randomly in the 2D plane. This detection probability increases with the dimension $n$ of the field $GF(2^{n})$.
  • $\bullet$
    Experimental results highlight the importance of choosing the field size carefully. For small fields such as $GF(2^{6})$, natural structural interactions between points can invalidate the probabilistic analysis. However, for larger fields such as $GF(2^{32})$ and beyond, the analysis holds reliably.
  • $\bullet$
    The proposed scheme can be naturally extended to support multiple secrets. We present two options: (1) using multiple lines, each encoding a distinct secret, and (2) using higher-degree polynomials instead of lines. In the latter case, reconstructing the embedded polynomial corresponds to solving the Maximum Co-Polynomial Points problem. However, existing algorithms for this task have complexities that grow with the polynomial’s degree.
Table 2 compares some multi-bit watermarking schemes with the one presented in this work.
Table 1: Some LLM-generated text multi-bit watermarking schemes
Authors Tools User Identification Integrity Complexity / Efficiency
Qu et al [18] RS Codes, Hash RS decoding
Jiang et al [11] TTP, Hash Efficient – needs TTP
This work Lagrange Interpolation, Hash, NFSR Solving MCP
The rest of the paper is structured as follows. Section 3 focuses on the quality and security properties required from a watermarking scheme. Section 4 describes the tools and building blocks used in the design of our watermarking schemes. Section 5 introduces our basic watermarking scheme along with algorithms for recovering watermarking bits. Section 6 presents an algebraic analysis of the scheme and a variant that is immune to such attacks. It also includes a probabilistic analysis evaluating the likelihood that an adversary can implant a fake point that would be accepted as genuine. Section 7 explores techniques for fast watermark extraction and double-resolution watermarking. Section 8 details our experimental results. Section 9 demonstrates how our watermarking scheme can be extended to support multiple secrets. Finally, Section 10 concludes the paper.
3 Quality and Security Requirements for LLM-text Watermarking
The main requirements for LLM-text watermarking ensure that the watermark is effective (i.e. easy to embed and extract) while preserving text quality. These include: imperceptibility, textual coherence and readability, scalability and efficiency, easy and reliable detection, and adaptability to different LLMs. Typical adversarial models for LLM-text watermarking assume that the owner of the text (either the LLM itself or its agent) embeds a watermark into the generated content. Once the watermarked text is published, an adversary with access to the text may attempt the following attacks:
  • $\bullet$
    Recovery of the watermark from the text to analyze or reverse-engineer the embedding method.
  • $\bullet$
    Removal of the watermark to evade detection and attribution.
  • $\bullet$
    Impersonation, where the adversary embeds a watermark to mimic a different AI model or author.
  • $\bullet$
    Modification of the watermark through insertion, repetition, substitution, or deletion, with the goal of making the changes undetectable while disrupting verification.
  • $\bullet$
    Collusion attack, where the adversary, having access to multiple watermarked text samples, combines them into a single text that passes watermark verification.
  • $\bullet$
    Extraction of secret keys by applying cryptanalysis techniques to break the watermarking scheme.
The required security features for LLM-text watermarking are enumerated below:
  • $\bullet$
    Resistance against adversarial editing – the watermark should remain detectable even if the text undergoes modifications such as paraphrasing, truncation, or reformatting.
  • $\bullet$
    Resilience to modifications, substitutions, and removal of watermarked text.
  • $\bullet$
    Authentication – the watermark should provide strong proof that the text was generated by a specific source (either by LLM or author who holds a secret key).
  • $\bullet$
    Low false positive (FP) and false negative (FN) rates – the detection method should minimize errors, ensuring that non-watermarked text is not falsely flagged and that genuine watermarks are not missed.
  • $\bullet$
    Security against collusion – the watermark should remain secure even if multiple watermarked texts are combined or analysed to disclose the secret elements (keys).
In this work, we focus on a specific security goal: enabling the owner of an LLM-generated text to prove that the text contains a substantial portion of their original content, from which a secret identity $K$ can be reliably recovered. This identity serves as a cryptographic witness of authorship or ownership. We further show how to extend our scheme so it provides integrity verification. Our approach is grounded in the use of Lagrange interpolation over points in a two-dimensional plane. Each point encodes a fragment of the embedded identity. To ensure unpredictability and prevent pattern leakage, we employ a pseudorandom bit generator (PRBG) to determine the $x$-coordinates of the interpolation points.
4 Components for Watermark Designs
4.1 Lagrange Interpolation over $GF(2^{n})$
Lagrange interpolation over a finite field $GF(2^{n})$ constructs a unique polynomial $f(x)\in GF(2^{n})[x]$ of degree at most $t-1$ that passes through a given set of $t$ distinct points $(x_{1},y_{1}),\dots,(x_{t},y_{t})$, where all $x_{i}\in GF(2^{n})$ and $x_{i}\neq x_{j}$ for $i\neq j$. The interpolating polynomial is given by:
$f(x)=\sum_{j=1}^{t}y_{j}\cdot\ell_{j}(x),$
where each $\ell_{j}(x)\in GF(2^{n})[x]$ is a Lagrange basis polynomial defined as:
$\ell_{j}(x)=\prod_{\begin{subarray}{c}1\leq m\leq t\\ m\neq j\end{subarray}}\frac{x-x_{m}}{x_{j}-x_{m}}.$
All arithmetic operations are performed in the field $GF(2^{n})$, including the inverses $(x_{j}-x_{m})^{-1}$. The basis polynomials satisfy $\ell_{j}(x_{i})=\delta_{ij}$, ensuring that $f(x_{i})=y_{i}$ for all $i$, where $\delta_{ij}$ denotes the Kronecker delta ($\delta_{ij}=1$ if $i=j$ and 0 otherwise). This form of interpolation is widely used in finite field cryptography, such as in Shamir’s secret sharing, error-correcting codes, and information dispersal schemes.
4.2 Pseudorandom Bit Generation
We need to generate $x$-coordinates at random. To achieve this, we rely on generators that produce bit sequences exhibiting random-like behaviour, even though they are deterministic algorithms initialised by a seed. These generators generally fall into three main categories: linear feedback shift registers (LFSRs), nonlinear feedback shift registers (NFSRs), and cryptographically strong pseudorandom bit generators (CSPBGs).
A linear feedback shift register is a shift register with linear feedback, typically using XOR, that generates deterministic sequences from an initial state. With a well-chosen feedback polynomial, an $m$-bit LFSR can produce a maximum-length sequence (m-sequence) cycling through all $12^{m}-$ 1 nonzero states. LFSRs are valued for their simplicity, speed, and efficient hardware implementation, making them popular in cryptography, error detection (e.g., CRC), random number generation, and spread-spectrum communication. Despite their advantages, LFSRs are vulnerable to algebraic attacks – observing just $m$ output bits can reveal the full state, and even nonlinear filtering does not fully mitigate this risk (see Courtois and Meier [7]).
A nonlinear feedback shift register is a generalization of the LFSR, in which the feedback function is highly nonlinear. This added complexity increases resistance to cryptanalytic techniques such as linear and algebraic attacks. NFSRs are widely used in cryptographic applications due to their strong security properties and efficient implementation in both software and hardware. A notable selection of NFSR-based designs was proposed during the CAESAR competition [3].
A cryptographically strong pseudorandom bit generator is a deterministic algorithm that produces bit sequences that are computationally indistinguishable from true randomness. Unlike general-purpose pseudorandom generators, CSPBGs are specifically designed to resist cryptographic attacks, ensuring that an adversary cannot feasibly predict future output even after observing a polynomial number of bits. Their security is typically based on hard mathematical problems, such as the shortest vector problem in lattices. While highly secure, CSPBGs are often less efficient due to the computational complexity of the underlying cryptographic assumptions.
To maintain efficient implementation, LFSRs are preferred, particularly when the LLM-generated text is short, limiting the adversary’s ability to gather sufficient information to compromise the watermarking scheme, or when security is not the primary concern. In scenarios where stronger security is required, NFSRs serve as a robust alternative to LFSRs, offering enhanced resistance to cryptanalysis. CSPBGs, however, are generally not recommended for this context due to their inherent trade-off: achieving high security typically relies on solving computationally intractable problems, such as operations over high-dimensional lattices, which are resource-intensive and slow.
4.3 Embedding Bits into Tokens
Large language models (LLMs) employ token vocabularies of varying sizes, with recent models like GPT-3 and Llama 2 using over 30,000 tokens, while Llama 3 supports up to 128,000. The trend toward larger vocabularies improves tokenisation and model performance but comes at the cost of increased computational demands. To embed watermarks in generated text, we build upon the method of partitioning the vocabulary into two disjoint subsets ${\cal V}_{0}$ and ${\cal V}_{1}$ representing binary marks 0 and 1, respectively. This concept extends the idea of “Green” and “Red” tokens introduced by Kirchenbauer et al. [12], which designate allowable and disallowed tokens.
Token partitioning can be done efficiently using pseudorandom bit generators seeded with a secret key or derived using cryptographic hashes. However, regenerating subsets frequently can become inefficient. Instead, a token’s membership in ${\cal V}_{0}$ or ${\cal V}_{1}$ can be computed on the fly using methods such as linear equations over token ID bits, encryption (e.g., AES in ECB mode), or cryptographic hashes (e.g., SHA3). These approaches aim to maintain balanced partitions and are especially effective in high-entropy text, where the natural variety of token usage allows for seamless watermarking without degrading the quality or fluency of the output.
Consider an example where we apply the SHA-3 cryptographic hash function to token IDs, optionally combined with a secret key $s_{k}$. In this case, we compute $h=(h_{1},\ldots,h_{n})=\text{SHA3}(s_{k},ID_{v})$. The watermark assigned to the token $v$ is then defined as $m=\bigoplus_{i=1}^{n}h_{i}$. A well-designed hash function ensures that the subsets ${\cal V}_{i}$; $i\in\{0,1\}$, are balanced with high probability, providing a fast and efficient solution.
5 Watermark based on Lagrange Interpolation
In many cases, the owner of an LLM-generated text may need to prove authorship by demonstrating that the embedded watermark enables recovery of their secret identity $K$. In other words, the extraction algorithm should be capable of retrieving this identity even if the text has been modified by an adversary. To ensure verifiable authorship, any sufficiently long fragment of the watermarked text should contain enough information to reveal the owner’s identity, serving as undeniable proof. We propose Lagrange interpolation as a convenient tool for recovering a secret polynomial $f(x)$, whose coefficients encode the identity $K$, from points in the two-dimensional plane with randomly generated $x$-coordinates. This guarantees that any sufficiently large portion of the text yields enough points to reconstruct $K$. The approach is resilient to partial alterations, making it robust in adversarial scenarios where the text may be tampered with to hide authorship.
As previously mentioned, we use Lagrange interpolation for polynomials of degree $t=2$. In this case, we work with a straight line $f(x)=a_{0}+a_{1}x\in GF(2^{n})[x]$, where the secret identity is encoded as $K=(a_{0}\|a_{1})\in GF^{2}(2^{n})$. Suppose our watermark consists of $N$ points lying on $f(x)$, implying that the number of tokens in the LLM text is $T=2nN$. An LFSR with secret seed $s_{k}$ is used to generate the coordinates $x_{i}$. For each $x_{i}$, we compute the pair $(x_{i},f(x_{i}))$, resulting in the watermark sequence:
$(x_{1},f(x_{1})),\ldots,(x_{N},f(x_{N})),$
where $x_{i}=LFSR(t_{i},s_{k})$ and $t_{i}$ denotes the clock instance. Any two such points are sufficient to reconstruct the secret $K$. Now assume the original LLM text $\cal T$ has been tampered with by an adversary, resulting in a modified version $\widetilde{\cal T}$. A verification algorithm extracts the watermark from $\widetilde{\cal T}$, possibly containing errors or missing points. Crucially, all correctly recovered points will still lie on the original line $f(x)=a_{0}+a_{1}x$. The verifier can accept $K=(a_{0}\|a_{1})$ as a valid candidate if it corresponds to the line with the most points recovered from $\widetilde{\cal T}$. If a recovered point is $(\tilde{x}_{i},\widetilde{f(x_{i})})$ instead of the correct $(x_{i},f(x_{i}))$, the probability that it still lies on the correct line is $\nicefrac{{1}}{{2^{n}}}$. This holds because for any incorrect $\tilde{x}_{i}$, there exists exactly one value $f(\tilde{x}_{i})$ that would make the point valid.
Fortunately, our adversary is unlikely to remove, replace, or insert single tokens, as such changes are easily detectable as typos. In other words, a single change (that is not a typo) typically generates a sequence of changes enforced by the LLM statistics (or the adversary intention). For this reason, the adversary is more likely to operate on larger sequences of tokens (such as sentences or paragraphs). This also means that the adversary tampers with longer sequences of tokens. A typical English language sentence consists of 10 to 20 tokens, while a paragraph contains between 150 and 200 tokens. Moreover, there is a trend in the development of new LLMs that are capable of processing more and more tokens.
Algorithm 5 outlines the process of watermark embedding using Lagrange interpolation. We depart from the conventional approach where the lists ${\cal V}_{0},{\cal V}_{1}$ are selected anew for each token, as this introduces unnecessary computational overhead. Assume we wish to embed the bit $i\in\{0,1\}$. First, the LLM is queried to produce logits. We then adjust the probabilities of the top tokens (based on the logits) that belong to ${\cal V}_{i}$ by checking whether the corresponding parity condition is satisfied. The parity check may be defined either as the XOR of the token ID bits or, alternatively, as a parity function applied to the output bits of a hash function $H(s_{k},ID_{\mbox{\it token}})$ – see Section 4.3. In this approach, the algorithm uses static sets ${\cal V}_{0}$ and ${\cal V}_{1}$ for embedding watermark bits across the text. While it is straightforward to use dynamic lists to enhance security, doing so increases the computational cost of watermark extraction.
Algorithm 1: Watermark Embedding Secret Identity $K$ 1:Input: Autoregressive language model LLM used to generate a text of length $T=2nN$, prompt sequence $PR$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, LFSR with $n$-bit state and straight line $f(x)=a_{0}+a_{1}x$ with the secret identity $K=(a_{0}\|a_{1})$, where computations are done in $GF(2^{n})$. 2:Output: Public LLM text $S_{1,1},\dots,S_{N,2n}$.   3:Initialize $LFSR(t_{0},s_{k})$ 4:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 5:for $i=1$ to $N$ do 6:     $x_{i}=(x_{i,1},x_{i,2},\ldots,x_{i,n})\leftarrow LFSR(t_{i},s_{k})$ 7:     $y_{i}=f(x_{i})=(y_{i,1},y_{i,2},\ldots,y_{i,n})$ 8:     for $j=1$ to $n$ do 9:         Compute logits: $v_{1}=\text{LLM}(PR,S_{1,1},\dots,S_{i,2(j-1)})$ 10:         Bias the logits: 11:         if $x_{i,j}=0$ then add bias to $\hat{v}_{1}[w]=v_{1}+\delta$ for all $w\in{\cal V}_{0}$ 12:         else add bias to $\hat{v}_{1}[w]=v_{1}+\delta$ for all $w\in{\cal V}_{1}$ 13:         end if 14:         Append $S_{i,2j-1}$ to $S$ by sampling from biased logits $\hat{v}_{1}$ 15:         Compute logits: $v_{2}=\text{LLM}(PR,S_{1,1},\dots,S_{i,2j-1})$ 16:         Bias the logits: 17:         if $y_{i,j}=0$ then add bias to $\hat{v}_{2}[w]=v_{2}+\delta$ for all $w\in{\cal V}_{0}$ 18:         else add bias to $\hat{v}_{2}[w]=v_{2}+\delta$ for all $w\in{\cal V}_{1}$ 19:         end if 20:         Append $S_{i,2j}$ to $S$ by sampling from biased logits $\hat{v}_{2}$ 21:     end for 22:end for 23:return $S$
The algorithm processes text block by block. For each block (lines 6,7), two watermark sequences are prepared. The first is a block generated by LFSR, which consists of an $x$-coordinate. The second block $y_{i}=f(x_{i})$. The loop from line 8 to 21 injects bits of pairs $(x_{i},f(x_{i}))$ into the LLM text.
Algorithm 2: Watermark Extraction 1:Input: LLM text $\widetilde{\cal T}=(\tilde{S}_{1,1},\dots,\tilde{S}_{N,2n})$, token set ${\cal V}$, $n$-bit secret key $s_{k}$, $n$-bit state LFSR, which generates $x$-coordinates of unknown $f(x)$, where points $(x_{i},f(x_{i}))$ are embedded into the text according to Algorithm 5; $i=1,\ldots,N$, computations are done in $GF(2^{n})$. 2:Output: The secret user identity $K=(a_{0}\|a_{1})$   3:Create an empty list of points ${\mathbb{P}}=\varnothing$ of $(x,f(x))$ 4:Initialize $LFSR(t_{0},s_{k})$ 5:Create the sequence ${\mathbb{L}}=(x_{1},\ldots,x_{N})$, where $x_{i}=LFSR(t_{i},s_{k})$ 6:Using secret key $s_{k}$, create two lists ${\cal V}_{0}$, ${\cal V}_{1}$, where ${\cal V}_{i}$ are tokens carrying watermark $i\in\{0,1\}$ (${\cal V}_{0}\cup{\cal V}_{1}={\cal V}$) 7:Extract watermark ${\mathbb{W}}=(w_{1},\ldots,w_{2nN})$ using ${\cal V}_{0}$ and ${\cal V}_{1}$ 8:for $i=1$ to $N$ do 9:     take $x_{i}$ 10:     while $x_{i}\neq(w_{j},\ldots,w_{j+n})$ do 11:         move $x_{i}$ along ${\mathbb{W}}$ bit by bit for $j=1,\ldots,n(2N-1)$ 12:     end while 13:     if $x_{i}=(w_{j},\ldots,w_{j+n})$ then 14:         ${\mathbb{P}}\leftarrow{\mathbb{P}}\cup(x_{i},(w_{j+n+1},\ldots,w_{j+2n}))$ 15:     end if 16:end for 17:return $\tilde{K}=(a_{0}\|a_{1})$ if the set ${\mathbb{P}}$ contains at least $\varepsilon\ell$ points on the straight line $f(x)=a_{0}+a_{1}x$
Algorithm 5 outlines the process of watermark extraction. The procedure starts by initializing the list $\mathbb{P}$ of points and computing the variables $x_{i}=x_{i}$ for $i=1,\dots,N$ using the LFSR. Two lists, ${\cal V}_{0}$ and ${\cal V}_{1}1$, are then created. Once these lists are prepared, the watermark ${\mathbb{W}}=(w_{1},\ldots,w_{2nN})$ is extracted. Each variable $x_{i}$ is processed sequentially, with the algorithm sliding $x_{i}$ bit by bit along ${\mathbb{W}}$ to search for a match. When a match is found, the next $n$ bits of ${\mathbb{W}}$ are assumed to be the suspected $f(x_{i})=(w_{j+n+1},\ldots,w_{j+2n})$, and the corresponding point is stored in the list $\mathbb{P}$. Processing each $x_{i}$ results in at most $2n(N-1)$ points, meaning that the set $\mathbb{P}$ has a cardinality of at most $2n(N-1)N$. If the text has been tampered with by an adversary, some points may no longer lie on the straight line $f(x)=a_{0}+a_{1}x$, and these outliers can be treated as random points. As long as enough points remain on the correct straight line, it can still be reliably determined, provided that adversarial distortions do not generate competing lines with a similar number of points. The primary challenge in this process is identifying the straight line containing the largest subset of points from the set, which is a well-known problem called the Maximum Collinear Points (MCP) problem.
5.1 Maximum Collinear Points Problem
The adversarial model is illustrated in Figure 1. In this model, the adversary has full control over the text and may apply arbitrary modifications, including insertions and deletions. However, extensive editing is often counterproductive, as it effectively amounts to rewriting the entire text from scratch. In practice, it is reasonable to assume that the adversary will aim to preserve as much of the original content as possible to minimise effort – after all, the motivation behind plagiarism is typically to save time by reusing someone else’s work. Interestingly, if the original watermark is available, the secret $K$ can be recovered directly using Lagrange interpolation. However, a verifier who only knows the key $s_{k}$ faces a more challenging task, as the extracted watermark may have been tampered with by an adversary. Effectively, the verifier deals with an instance of the MCP problem.
Original LLM text Original Watermark Adversary Modified LLM text Modified Watermark Lagrange Interpolation Secret $K$ MCP algorithm Secret $\tilde{K}$
Figure 1: Adversarial model
The MCP problem is defined as follows:
Instance: A set $\mathbb{S}$ of $N$ points in the 2D plane, where each point $p_{i}=(x_{i},y_{i})$; $i=1,2,\ldots,N$.
Question: What is the largest subset ${\mathbb{S}}^{\prime}\subset{\mathbb{S}}$ of points that are collinear (i.e., lie on a single straight line)?
There are many efficient algorithms that solve the problem. The basic ones are as follows:
  • $\bullet$
    Brute force approach — the algorithm checks all triplets for collinearity. Its time and space complexities are ${\cal O}(N^{3})$ and ${\cal O}(1)$, respectively (see [6]).
  • $\bullet$
    Sorting-based approach — the algorithm takes each point as a reference and computes slopes between the reference point and the other points. The slopes are sorted to group collinear points. Finally, it counts the maximum number of collinear points for each reference point. Its time and space complexities are ${\cal O}(N^{2}\log{N})$ and ${\cal O}(N)$, respectively (see [2]).
  • $\bullet$
    Hashing-based approach — the algorithm is similar to the sorting-based one, except that instead of sorting, it applies hashing to store slopes, reducing the time complexity to ${\cal O}(N^{2})$ (see [6]).
  • $\bullet$
    Convex-hull approach — the algorithm uses a convex hull (such as Graham’s scan or Andrew’s monotone chain) to preprocess points and then adds points, checking them for collinearity. Its time and space complexities are ${\cal O}(N^{2})$ and ${\cal O}(N)$, respectively (see [17]).
The algorithms presented above show that the MCP problem can be efficiently solved for the number of points $N<10^{5}$. Table 2 compares the algorithms.
Algorithm Time Complexity Space Complexity Approach
Brute force [6] ${\cal O}(N^{3})$ ${\cal O}(1)$ Check triplets
Sorting-based [2] ${\cal O}(N^{2}\log N)$ ${\cal O}(N)$ Sort slopes
Hashing-based [6] ${\cal O}(N^{2})$ ${\cal O}(N)$ Hashing for slopes
Convex hull [17] ${\cal O}(N^{2})$ ${\cal O}(N)$ Geometric hull
Table 2: Comparison of algorithms for the Maximum Collinear Points Problem
Let us consider the Hashing-based algorithm (see Algorithm LABEL:alg:hash_MCP). It determines the largest subset of collinear points among $N$ given points in a 2D plane. The algorithm iterates through each point, treating it as a reference. For every other point, it computes the slope relative to the reference and stores the slope count in a hash map. This hash map allows for quick lookups and updates, ensuring that collinear points are efficiently counted. The maximum count of any slope within the hash map, plus the reference point itself, gives the largest number of collinear points for that reference. By iterating through all points, the algorithm determines the global maximum. Note that vertical lines can be ignored as they are not parts of secret sharing.
Algorithm 3: Hashing-based MCP (Maximum Collinear Points) 1:Input: A set of $N$ points ${\mathbb{S}}=\{p_{1},p_{2},\dots,p_{N}\}$ in $GF(2^{n})\times GF(2^{n})$ 2:Output: Maximum number of collinear points   3:maxCount $\leftarrow 1$ 4:for each point $p_{i}\in{\mathbb{S}}$ do 5:     Create an empty hash map $H$ 6:     for each point $p_{j}\in{\mathbb{S}}$ such that $p_{j}\neq p_{i}$ do 7:         Compute slope $s=\frac{y_{j}-y_{i}}{x_{j}-x_{i}}$triangleright Handle vertical lines separately 8:         Increment $H[s]$ by 1 9:     end for 10:     maxCount $\leftarrow\max($maxCount, $\max_{s\in H}(H[s]+1))$ 11:end for 12:return maxCount
6 Security Evaluation
6.1 Algebraic Analysis
By now, the reader should notice that the proposed watermarking scheme relies on two linear components: the LFSR and the function $f(x)$. The only nonlinear element arises from the division of the token set into ${\cal V}_{0}$ and ${\cal V}_{1}$, which occurs when cryptographic hashing or encryption is employed, as discussed in Section 4.3. It is important to observe that Algorithm 5 uses fixed sets ${\cal V}_{0}$ and ${\cal V}_{1}$ throughout the entire embedding process. This design choice introduces a potential vulnerability, as it enables the adversary to mount the following attack.
Algebraic Attack on Watermarking Scheme
  1. 1.
    Target the most frequent tokens that appear in the attacked LLM-generated text. Determine $n$ positions, where the top most frequent tokens occur.
  2. 2.
    For each position and its token, write a binary linear relation of the form:
    $\bigoplus_{i=1}^{n}\delta_{i}s_{i,k}=w_{j},$
    where $s_{k}=(s_{1,k},\ldots,s_{n,k})$ is the LFSR seed, and $w_{j}$ is an unknown bit. The value $w_{j}$ is either 00 or $1$, depending on whether the token belongs to ${\cal V}_{0}$ or ${\cal V}_{1}$.
  3. 3.
    Collect the resulting system of linear equations and solve them for guessed values of $w_{j}$. If the adversary observes the same token in $n$ positions, they can solve the system for both $w_{j}=0$ and $w_{j}=1$, and be confident that one solution is correct.
  4. 4.
    Verify the candidate solution by substituting the computed $\tilde{s}_{k}$ into the LFSR and checking whether the generated sequence matches the observed token positions. This reveals how watermarking bits are assigned to tokens, assuming the LFSR governs watermark generation.
  5. 5.
    Using the identified tokens, assemble $n$ binary relations associated with the function $f(x)$ and recover the secret identity $K$. This is feasible when some tokens carrying values $f(x_{i})$ have been identified via their membership in ${\cal V}_{0}$ or ${\cal V}_{1}$. As a result, the watermarking scheme can be fully compromised.
Algebraic attacks can be mitigated using simple yet effective countermeasures, as outlined below.
  • $\bullet$
    Instead of using an LFSR, it is advisable to employ a cryptographically secure NFSR, such as those proposed in the CAESAR competition [3]. While an adversary may still attempt to construct algebraic relations for the most frequent tokens, each guess of watermark bits leads to highly nonlinear equations, rendering the system intractable. For security-critical applications, it is recommended to use provably secure cryptographic pseudorandom bit generators.
  • $\bullet$
    As an alternative – or in addition – to the above, the fixed sets ${\cal V}_{0}$ and ${\cal V}_{1}$ can be reshuffled periodically. For instance, the token set can be repartitioned every $n$-bit or $2n$-bit block. Furthermore, different partitions can be employed for the $x_{i}$ blocks generated by the LFSR and for the $f(x_{i})$ blocks. In the most extreme approach, ${\cal V}_{0}$ and ${\cal V}_{1}$ can be reshuffled for every individual token. However, this significantly increases computational overhead, especially during watermark extraction.
Watermarking Scheme Immune to Algebraic Attacks Given the watermarking scheme with embedding and extraction described in Algorithms 5 and 5, the following modifications make the scheme immune to algebraic attacks: 1. Apply a nonlinear feedback shift register (NFSR) instead of an LFSR to prevent recovery of the seed $s_{k}$. 2. Replace fixed token space partitions with dynamic ones, depending on the type of token: $\bullet$ For tokens corresponding to $x_{i}$ generated by NFSR, use a fixed partition of ${\cal V}$. For instance, assign a token $v$ with identity $ID_{v}$ a watermark bit $m=\bigoplus_{j=1}^{n}h_{j}$, where $(h_{1},\ldots,h_{n})=\mathrm{SHA3}(s_{k},ID_{v})$. $\bullet$ For $y_{i}=f(x_{i})$, create a new partition for each $i$, breaking the linear dependencies induced by $f(x)$. For example, define the token partition using a hash $H(s_{k},x_{i})$. A specific instance might compute $(h_{1},\ldots,h_{n})=\mathrm{SHA3}(s_{k},x_{i},ID_{v})$. 3. This design enables efficient watermark extraction with only a slight increase in complexity. The verifier first applies the partitioning for tokens $x_{i}$ generated by NFSR. Once all $x_{i}$ are identified, the verifier extracts the watermark values $y_{i}$ from the corresponding $n$-token blocks.
6.2 Probability of Watermark Extraction
Let us consider the extent to which our adversary must tamper with our watermarking to destroy the watermark. In other words, we aim to determine the threshold at which the watermark’s owner can still prove that the LLM-generated text has been watermarked by them or extract the owner’s identity $K$. We can extract $N=F+R$ points, where $F$ belong to the straight line $f(x)$ and $R$ have been tampered with by the adversary. It is reasonable to assume that the $R$ points are randomly scattered on the plane $GF(2^{n})\times GF(2^{n})$. Clearly, the $K$ extraction is successful if any subset of random points forming a straight line is no larger than $F$.
First, consider the following trivial facts about straight lines on the $GF(2^{n})\times GF(2^{n})$ plane:
  • $\bullet$
    Any two points uniquely determine a straight line. This means that there are $2^{2n}$ straight lines.
  • $\bullet$
    Any two straight lines can either intersect at a single point or be parallel with no points in common.
  • $\bullet$
    For each point, there are $2^{n}$ straight lines that contain that point.
The above properties show that for the number $R$ of random points larger than $2^{n}$, the correlation among the straight lines grows. Consequently, finding the precise probability distribution of the number of random points that belong to different straight lines is a challenging problem.
Theorem 6.1.
Given a plane $GF(2^{n})\times GF(2^{n})$, $R$ random points and the assumption that $R\ll 2^{n}$, then the probability that there is a straight line that contains $k$ points is approximated by
$\binom{R}{k}\left(\frac{1}{2^{n}}\right)^{k-2}\left(1-\frac{1}{2^{n}}\right)^{R-k},$
where $k>2$.
Proof.
We start tossing $R$ points. The first two tosses give a unique straight line. The event that the tosses produce the same point occurs with negligible probability so we ignore it. Now, we expect that among the $(R-2)$ other tosses, $(k-2)$ belong to the line (with probability $\nicefrac{{1}}{{2^{n}}}$) and $(R-k)$ fall outside the line (with probability $(1-\nicefrac{{1}}{{2^{n}}})$). The binomial coefficient provides the number of possible choices of $k$ out of $R$ possibilities. $\Box$
Theorem 6.1 provides us with a tool to evaluate the probability that Algorithm 5 extracts the identity $K$ correctly. The algorithm fails to give the true $K$ if there is a subset of $F$ or more random points (whose total number is $R$) sitting on the same straight line. The probability of this event is
$P(\mbox{fail})=\sum_{k=F}^{R}\binom{R}{k}\left(\frac{1}{2^{n}}\right)^{k-2}\left(1-\frac{1}{2^{n}}\right)^{R-k}$
Example 1.
Consider a toy example, where calculations are done in $GF(2^{8})$, i.e. $n=8$. We have $F=4$ points lying on a straight line and $R=16$ random points. The probability that four random points belong to a straight line is
$\binom{16}{4}\left(\frac{1}{64}\right)^{2}\left(\frac{63}{64}\right)^{12}\approx 0.37$
It is easy to check that probabilities that five and more points lie on a straight line vanish exponentially. Consequently, the probability of correct recovery of $K$ is $\approx 0.62$. The probability of the correct $K$ recovery improves significantly if $F=5$ and $R=15$ as it increases to $\approx 0.99$.
7 Extensions and Variants
7.1 Simplified Watermark Extraction
Algorithm 5 recovers watermarks by identifying the values $\mathbb{L}=(x_{1},\ldots,x_{N})$, which serve as the $x$-coordinates of the points. After locating an $x_{i}$ that was generated by LFSR($s_{k}$), the subsequent block of $n$ bits is extracted and interpreted as $f(x)$. In other words, our MCP algorithm operates on at most $N$ points, as those with incorrect $x_{j}$ values are ignored. This selective approach is particularly effective when dealing with long texts. However, for shorter texts, we may be inclined to run the MCP algorithm directly – without filtering out points with invalid $x$-coordinates. The alternative idea is to consider $12n(N-1)+$ 1 points, which are generated by sliding a $2n$-bit window bit by bit across the entire $2nN$-bit watermark. Our experiments confirmed expectations: apart from the $N$ collinear original points, the additional points produced through bit shifts behave like random points and do not interfere with the collinear structure. However, special attention must be paid to small Galois fields such as $GF(2^{6})$. In such cases, the random selection of $x$-coordinates may result in duplicate points, and the quality of randomness from the sliding-window operation may be questionable. Table 3 shows results, when $x$-coordinates are selected (pseudo)-randomly.
Arithmetics # Collinear Points # Points by Shifting
5 49
$GF(2^{6})$ 10 109
20 (18) 229
$GF(2^{12})$ 10 109
80 1897
$GF(2^{32})$ 40 2497
100 6337
Table 3: Collinear points when the MCP algorithm is fed with points generated by shifting
Note that the red entry for $GF(2^{6})$ indicates that our MCP algorithm returned 18, which is clearly incorrect. Since we used pseudorandom generation, there is a high likelihood that some of the 20 random points are duplicates, given that there are only 64 possible $x$-coordinates to choose from.
7.2 Double-Resolution Watermarking
We embed watermarks by defining a sequence of points $(x_{i},f(x_{i}))\in GF^{2}(2^{n})$; $i=1,\ldots,N$. When using our simplified extraction method, the MCP algorithm must process $12n(N-1)+$ 1 points. However, the number of points generated by shifting a $2n$-bit window can be reduced to $n(N-1)+1$ by overlapping point coordinates. In other words, instead of using the coordinates $x_{1},x_{2},\ldots,x_{N}$ generated by an LFSR, we can let $f(x)$ recursively generate the coordinates as $x,f(x),f^{2}(x),\ldots,f^{N-1}(x)$. Consequently, the sequence of points becomes
($\underset{P_{1}}{\underbrace{(x,f(x))}},\ \underset{P_{2}}{\underbrace{(f(x),f^{2}(x))}},\ \ldots,\ \underset{P_{N}}{\underbrace{(f^{N-1}(x),f^{N}(x))}}$
The sequence can be compressed into
$(x,f(x),f^{2}(x),\ldots,f^{N-1}(x),f^{N}(x))$
It still preserves the $N$ collinear points, while reducing the number of “random” points introduced by the shifting operation. There are, however, some caveats to consider. First, repeated application of the function $f(x)$ leads to a cycle. If this cycle is shorter than $N$, then a new initial $x$ must be selected. Second, and more importantly, this design may have security implications that must be carefully addressed.
8 Experiments
We conducted experiments to analyze the maximum number of collinear points that can be generated by an adversary who randomly selects points on a 2D plane. The adversary’s goal is to deceive the verifier by creating a dense enough cluster of points that form a fake straight line, which the verifier might mistake for a genuine watermark. These experiments were carried out using the PARI software package - see Table 4.
Random Points # GF($2^{6}$) GF($2^{8}$) GF($2^{12}$) GF($2^{16}$) GF($2^{32}$)
8 3 2 2 2 2
16 3 3 2 2 2
32 5 3 2 2 2
64 6 4 3 2 2
128 6 5 3 2 2
256 10 5 3 2 2
512 14 8 3 2 2
1024 20 11 4 2 2
Table 4: Number of collinear points detected in random sets of varying sizes across different Galois fields
The results confirm our intuition: in smaller fields, the number of collinear points tends to grow, but not significantly. In contrast, for larger fields such as $GF(2^{32})$ and beyond, the maximum number of collinear points observed is typically just two – indicating that the points are largely scattered across the plane, forming only trivial two-point lines. Under these conditions, the presence of three or more collinear watermarking points is a strong indicator of intentional embedding. Moreover, the sender can embed a larger number of such points to make the watermark more resilient against adversarial interference.
Given a short text, such as a paragraph consisting of 300 tokens, we can embed a 300-bit watermark. In practice, a key challenge lies in selecting the appropriate dimension of the Galois field used for the computation. If we choose $GF(2^{6})$, each point occupies 12 bits. This means the text can accommodate 25 watermark points. According to Table LABEL:tab:mcp, only 5 of these points need to survive for the correct watermark line to be recovered, while the remaining 20 points may be replaced by fake or random ones without affecting recoverability. If we instead use $GF(2^{8})$, the total number of watermark points decreases to 18. In this case, as shown in Table LABEL:tab:mcp, just 4 genuine points are enough to reconstruct the correct line, while the remaining 14 random points form lines with a maximum of only 3 points. With $GF(2^{12})$, the number of embedable points drops further to 12. Still, only 3 genuine watermark points are needed to reconstruct the original line, as 9 random points only form trivial two-point lines. A broader discussion considering different text lengths – such as a paragraph, a page, and a 10-page report – is provided in Table 5. It is evident that for longer documents, using Galois fields of higher dimension is advantageous, as the appearance of a three-point line becomes strong evidence of the watermark’s presence and authenticity.
Document Number of tokens Arithmetics Total number Number of Number of
(bits) GF of points watermark points fake points
$GF(2^{6})$ 25 5 20
Paragraph 300 $GF(2^{8})$ 18 4 14
$GF(2^{12})$ 12 3 9
$GF(2^{6})$ 58 7 51
Page 700 $GF(2^{8})$ 43 5 38
$GF(2^{12})$ 29 3 26
$GF(2^{6})$ 583 20 563
Report 7,000 $GF(2^{8})$ 437 5 432
(10 pages) $GF(2^{12})$ 291 4 287
$GF(2^{16})$ 218 3 215
Table 5: Watermarking points for different documents and arithmetics
9 Muli-Secret Watermarking
9.1 Multiple Lines
Using a single line $f(x)=ax+b\in GF(2^{n})$, it is possible to obtain the secret $K=(a\|b)\in\{0,1\}^{2n}$. Further expansion of the secret is possible by embedding $t$ distinct lines instead of just one. In this case, our MCP algorithm requires a minor adjustment to track the top $t$ straight lines that yield the highest point counts. This results in a collection of lines $\{f_{i}(x)=a_{i}x+b_{i}\;|\;i=1,\ldots,t\}$ and their corresponding secrets $\{K_{i}=(a_{i}\|b_{i})\;|\;i=1,\ldots,t\}$. To form a single $2tn$-bit secret, we must determine an appropriate order for concatenation. A simple method proceeds as follows:
  • $\bullet$
    For each secret $K_{i}$, where $i=1,\ldots,t$, compute its digest: $d_{i}=H(K_{i})$, where $H()$ is a cryptographic hashing.
  • $\bullet$
    Sort the digests $\{d_{1},\ldots,d_{t}\}$ in increasing order and reorder the corresponding secrets $K_{i}$ accordingly.
  • $\bullet$
    Concatenate the reordered secrets into a single sequence.
9.2 High Degree Polynomials
An obvious generalisation is a choice of higher degree polynomials for our $f(x)$ instead of straight lines. This means that we deal with the polynomial over $GF(2^{n})$ of the degree $t-1$ of the following form
$f(x)=a_{0}+a_{1}x+,\ldots,a_{t-1}x^{t-1}\mbox{ where }t\geq 3$
Watermarks are points $\{(x_{i},f(x_{i})\;|\;i=1,\ldots,N\}$. Clearly, to determine the polynomial $f(x)$, it is enough to know $t$ points. We can apply the well-know Lagrange Interpolation to reconstruct $f(x)$. The secret $K=(a_{0}\;\|\;a_{1}\;\|\;\ldots\;\|\;a_{t-1}\}$. The problem of recovery, however, becomes interesting when some points are corrupted by our adversary. The verifier in this case faces a task to identify a polynomial of degree $(t-1)$, which contains the maximum number of points (assuming that the adversary corrupted a relatively small portion of points). This is to say that we face the Maximum Co-Polynomial Points (MCPP) problem.
The MCPP problem is defined as follows:
Instance: A set $\mathbb{S}$ of $N$ points in the 2D plane, where each point $p_{i}=(x_{i},y_{i})$; $i=1,2,\ldots,N$.
Question: What is the largest subset ${\mathbb{S}}^{\prime}\subset{\mathbb{S}}$ of points that lie on a single univariate polynomial $f(x)$ of degree at most $(t-1)$?
Some basic algorithms for solving the problem are given below.
  • $\bullet$
    Brute force polynomial interpolation — the algorithm iterates over all $\binom{N}{t}$ subsets of $t$ points, interpolates a $(t-1)$-degree polynomial and counts how many other points lie on the curve $f(x)$. Its complexity is ${\cal O}(N^{t+1})$, which quickly becomes impractical for large $N$ and $t$ [4].
  • $\bullet$
    Random Sampling + Verification – the algorithm selects many subsets of $t$ points, interpolates a polynomial for the given subset and verifies how many points lie on it. Its complexity is ${\cal O}(T(t^{2}+N)$, where $T$ is the number of trials.
  • $\bullet$
    Hough Transform for Polynomials – the algorithm uses the classic Hough Transform (used in computer vision for line detection) to parametrise $t$-degree polynomials. Each point votes for all curves passing through it – curves with most votes are detected. Its complexity ${\cal O}(Nq^{t-1})$, where $q$ is the space size parameter [1].
  • $\bullet$
    Gröbner Basis – the algorithm encodes the point constraints as a system of polynomial equations and use Gröbner bases to simplify and solve. Its complexity is exponential in $t$ although often better in practice [20].
Finding a polynomial $f(x)$ of degree $(t-1)$ from $t$ points becomes increasingly challenging as $t$ grows. However, our scenario is slightly different. If the watermarks remain unaltered by the adversary, any collection of $t$ such points will correctly determine the polynomial. In contrast, if the text has been modified, the points extracted from the tampered sections will not lie on the original polynomial $f(x)$. The verifier must distinguish between two types of points: good ones that lie on $f(x)$ and bad ones, which are scattered randomly in the plane $GF^{2}(2^{n})$. The following theorem evaluates the probability of success when the verifier randomly selects $t$ points, hoping they all lie on $f(x)$.
Theorem 9.1.
Given a collection ${\mathbb{P}}$ of all watermarks/points recovered from LLM-generated text, where ${\mathbb{P}}={\mathbb{F}}\cup{\mathbb{A}}$, and ${\mathbb{F}}$ are points lying on $f(x)$ while ${\mathbb{A}}$ consists of random points introduced by the adversary. Suppose the verifier selects $t$ points at random, computes $\widetilde{f(x)}$, and checks whether other points are consistent with it. Then, the expected number of trials required to succeed (i.e., to find at least one set of $t$ points entirely from ${\mathbb{F}}$) is
$\left(\frac{\ell}{f}\right)^{t},$
where $\ell=\#{\mathbb{P}}$ and $f=\#{\mathbb{F}}$.
Proof.
The probability of success in a single trial is $p=(\nicefrac{{f}}{{\ell}})^{t}$. If a trial fails, the next has probability $(1-p)p$ of being successful. More generally, the probability that the $k$-th trial is the first success is $(1-p)^{k-1}p$. This defines the geometric probability distribution $P(X=k)=(1-p)^{k-1}p$, whose expected value is $\nicefrac{{1}}{{p}}$. $\Box$
Example 2.
Let $\ell=1000$, and let the number of good points be $f=200$. If the watermark is defined by a polynomial of degree $t=3$, then the expected number of trials is $(\nicefrac{{1000}}{{200}})^{4}=625$.
10 Conclusions
This work introduces a novel watermarking scheme for LLM-generated text that enables the recovery of a secret author identity through geometric encoding. The core idea is to represent the secret as a straight line $f(x)$ over a finite field, and embed a sequence of points $(x_{i},f(x_{i}))$ lying on this line into the LLM text. These points serve as the watermark, recoverable even after partial editing or adversarial manipulation. We demonstrate that the scheme is lightweight, easy to implement, and well-suited for applications with moderate security requirements. For security-critical scenarios, we enhance the robustness of the scheme by generating the $x_{i}$ coordinates using cryptographically secure pseudorandom number generators, and by introducing a variable partitioning of token sets for each $n$-token block. These modifications increase resistance to algebraic and probabilistic attacks.
The scheme is naturally extendable to support multiple secrets. One straightforward approach is to embed multiple lines, each corresponding to an individual secret. While effective, this method increases the probability that an adversary may generate a fake point that aligns with one of the embedded lines, potentially resulting in false positives. Moreover, if the goal is to produce a single unified secret, a method for determining the correct order of concatenation among the individual secrets must be established. To address this, we propose the use of higher-degree polynomials instead of lines. In this approach, the secret is encoded as the coefficients of a polynomial $f(x)$ and watermark points are sampled from this polynomial. This not only reduces the adversary’s chances of fabricating a point that lies on the curve—since a random point lies on a polynomial of the degree $t$ with probability only $\nicefrac{{1}}{{2^{n}}}$ – but also enables straightforward concatenation of the coefficients into a single, longer secret. The trade-off, however, is an increase in computational complexity: reconstructing a polynomial of the degree $t$ requires solving a more resource-intensive interpolation problem, especially as $t$ or the field size grows.
Overall, the proposed watermarking framework offers a flexible balance between efficiency and security, and opens the door to new watermarking paradigms based on algebraic structures. Future work may focus on optimising multi-secret recovery, enhancing resilience against editing attacks, and exploring hybrid schemes combining lines and polynomials to balance robustness and performance.
ACKNOWLEDGMENT
This research was funded in part by National Science Centre, Poland, grant 2023/49/B/ST6/02580. For the purpose of Open Access, the authors have applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission.